# Quantum noise spectroscopy of superconducting critical dynamics and vortex fluctuations in a high-temperature cuprate

### v1

@@ Abstract
To quantitatively analyze the observed magnetic noise, we extract the BSCCO-induced decay rate, $\Gamma_{1}(T)$, by subtracting the intrinsic relaxation rate — measured from NV centers located far from the BSCCO sample — from the total relaxation rate of NV centers situated beneath the BSCCO [Fig. 3(d)]. At high temperatures, $T>T_{c}$, $\Gamma_{1}(T)$ is finite due to Johnson-Nyquist noise in the metallic phase kolkowitz2015probing ; agarwal2017magnetic ; hsieh2019imaging . At very low temperatures, $T\ll T_{c}$, the magnetic noise is significantly suppressed in the superconductor. Crucially, the form of this suppression offers insights into the nature of the pairing symmetry. While in conventional s-wave superconductors $\Gamma_{1}(T)$ is exponentially suppressed at low temperatures due to a fully gapped Bogoliubov quasiparticle spectrum  bardeen1957theory ; bardeen1957microscopic ; chatterjee2022single ; dolgirev2022characterizing ; kelly2024superconductivity ; li2024observation , the presence of nodal quasiparticles in d-wave layered superconductors implies a distinct power-law suppression of $\Gamma_{1}(T)\propto T^{2}$ dolgirev2022characterizing ; chatterjee2022single . Indeed, this expectation is consistent with our observations [Fig. 3(d), inset], corroborating the nodal d-wave nature of superconductivity in BSCCO keimer2015quantum .
@@
The most striking feature of Fig. 3(d) is the sharp and symmetric peak in $\Gamma_{1}$ near the critical temperatures $T_{c}$. This peak arises from strong fluctuations in the superconducting order parameter, resulting in enhanced current fluctuations that act as source of magnetic noise at the metal-superconductor phase transition. Crucially, this enhancement is observed across the entire BSCCO sample, as emphasized by the spatial line cut near $T_{c}$, Fig. 3(c). Interestingly, near $T_{c}$, $\Gamma_{1}(T)$ begins to show marked deviations from mean-field expectations for d-wave superconductors. In particular, mean-field theory predicts a divergence of $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$ up to logarithmic corrections dolgirev2022characterizing . In contrast, the experimentally measured $\Gamma_{1}$ at criticality is distinctly more singular, with $x\approx 1$ [Fig. 3(e)]. This deviation arises because BCS theory neglects both amplitude and phase fluctuations of the superconducting order parameter $\psi$, which play a significant role near the critical point emery1995importance ; maki1968critical . These effects are particularly important for thin-film superconductors where fluctuations are expected to be stronger tinkham . Consequently, quantitative characterizing the divergent time-scale associated with dissipative dynamics of the superconducting order parameter near $T_{c}$ remains an important open question, which we turn to next.
@@
To properly account for low-energy critical fluctuations, we consider the following phenomenological Langevin dynamics of the superconducting order parameter $\psi(\mathbf{r},t)$.
@@
where $F[\psi]$ is the Ginzburg-Landau free energy functional landau1980statistical ; cyrot1973ginzburg ; schuller2006time ; tinkham , $\eta$ is local white noise that captures the effect of coarse-graining, and $\gamma$ sets the rate of relaxation towards equilibrium hohenberg1977theory . Crucially, by setting $r(T)=a(T-T_{c})$ with $a>0$, we can analyze fluctuations of the order parameter about the minima of the free energy: $\langle\psi\rangle=0$ for $T>T_{c}$ (metallic) and $\langle\psi\rangle=\psi_{0}$ for $T<T_{c}$ (superconducting). Such fluctuations in turn leads to enhanced current-induced magnetic noise at the NV frequency $D_{\mathrm{gs}}$, enhancing the relaxation rate $\Gamma_{1}(T)$ near $T_{c}$. This enhancement of the noise spectral weight at low frequency is a consequence of critical slowing down of order parameter fluctuations: for weak-coupling superconductors it is quantified by the diverging timescale $\tau_{\rm GL}=8\hbar/(k_{B}|T-T_{c}|)$ larkin2005theory ; schmid1966time near $T_{c}$.
We find that our theoretical model (see Methods) can faithfully reproduce both the magnitude of the magnetic noise, and the divergence of $\Gamma_{1}(T)$ near $T_{c}$ with the experimentally observed power-law $x=1$ on both sides of the transition [Fig. 3(e)]. In addition, by fitting $\Gamma_{1}(T)$ with our model, we extract a quantitative time-scale for critical slowing down $\tau_{0}^{\rm fitted,M}\approx 1.2\,\tau_{\rm GL}$ on the metallic side, and $\tau_{0}^{\rm fitted,SC}\approx 0.2\,\tau_{\rm GL}$ on the superconducting side. Remarkably, our results indicate that the decay rates of the order parameter near criticality are comparable to the analytical weak-coupling results $\tau_{0}^{\rm wc,M}=\tau_{\rm GL}$ larkin2005theory on the metallic side and $\tau_{0}^{\rm wc,SC}=0.5\,\tau_{\rm GL}$ on the superconducting side schmid1966time ; KimDiessel , even though high $T_{c}$ cuprates like BSCCO are expected to be strong-coupling superconductors lee2006doping .
Further, our modeling of critical noise also allows us to extract the static critical exponent $\nu$ that determines the divergence of the order parameter correlation length $\xi(T)\propto|T-T_{c}|^{-\nu}$, and the dynamical critical exponent $z$ that relates the correlation time $\tau_{0}$ and correlation length as $\tau_{0}(T)\propto\xi^{z}$. The close agreement of the experimental observations with Langevin dynamics of $\psi$ implies that $\nu=1/2$, and $\tau_{0}(T)\propto|T-T_{c}|^{-1}\propto\xi^{2}$ implies $z=2$, consistent with expectations for a non-conserved order parameter hohenberg1977theory .
@@
Figure 4: In-field criticality and vortex dynamics. (a) BSCCO-induced relaxation rate, $\Gamma_{1}(T)$ (normalized), as a function of temperature at $H=0,40,200$ G. Inset: the measured peak magnitude of $\Gamma_{1}$ as a function of applied field, which is consistent with a linear dependence on the field strength $H$ (dashed line). (b) BSCCO-induced decoherence rate, $\Gamma_{2}(T)$, as a function of temperature at $H=55,120,200$ G. Inset: The measured $T^{\mathrm{XY8}}_{2}$ decoherence timescale as a function of temperature for NVs under and away from BSCCO at $H=200$ G. (c) Schematic of the noise spectrum arising from different dynamical processes in BSCCO. The shaded area marks the most sensitive frequencies of NV relaxation ($\Gamma_{1}$) spectroscopy and decoherence ($\Gamma_{2}$) spectroscopy. Since the NV’s intrinsic decoherence rate far exceeds its relaxation rate, $\Gamma_{2}$ is less sensitive to weak magnetic noise than $\Gamma_{1}$. (d) Left: a trapped vortex imaged in BSCCO under $H=40$ G at 50 K. Right: the measured relaxation time $T_{1}$ and decoherence time $T^{\mathrm{XY8}}_{2}$ of NV centers at different spatial positions across the vortex.
@@
In-field criticality and vortex dynamics — Next we explore how an out-of-plane magnetic field impacts critical fluctuations. Specifically, we apply a small bias magnetic field and perform $T_{1}$ noise spectroscopy at two distinct fields, $H=40$ G and $H=200$ G. Intriguingly, compared with the zero-field $\Gamma_{1}$, two distinct new features emerge [Fig. 4(a)]: (i) The peak in the measured noise, $\Gamma_{1}$, is shifted towards a slightly lower temperature. (ii) While the noise signal approaching criticality from the metallic side remains relatively sharp, from the superconducting side, it increases much more gradually and smoothly towards criticality, resulting in a broad and asymmetric spectrum. The former is expected, as a perpendicular field hinders phase coherence and lowers $T_{c}$. The latter can be attributed to a significantly larger fluctuation regime, dominated by the two-dimensional XY critical behavior Beasley ; halperin1979resistive ; curtis2024probing . Below $T_{c}$, an intermediate vortex liquid phase blatter1994vortices ; fisherfisherhuse ; Nelson_VL is expected to form. In this phase, the motion of the superconducting vortices, which trap magnetic flux quanta, can generate strongly fluctuating magnetic fields at the NV, resulting in enhanced noise.
@@
Further evidence for this scenario can be gleaned from the scaling of the peak magnitude of $\Gamma_{1}$ with the magnetic field $H$ [inset, Fig. 4(a)], where the density of vortices $n_{v}$ increases with increasing field strength. Within a simple model of diffusive vortex motion (see Methods), the local magnetic noise, measured via $\Gamma_{1}(H)$, scales as $D_{v}^{-1}$ where $D_{v}\sim v^{2}\tau_{v}$ is the vortex diffusion constant. We expect the vortex scattering time $\tau_{v}$ to scale inversely proportional to the vortex density $n_{v}$, and the mean-square speed $\langle v^{2}\rangle$ is primarily determined by thermal energy. Consequently, the measured magnetic noise scales as $\Gamma_{1}(H)\propto n_{v}\propto H$, in line with our observation [inset, Fig. 4(a)]. As the temperature decreases further, the vortices either freeze into a lattice or get pinned by defects, resulting in a decrease of $\Gamma_{1}(T)$. This conclusion is consistent with our measurement of noise over a line-cut on a large pinned vortex at low temperatures shown in Fig. 4(d) — $T_{1}$ does not change appreciably across the vortex, indicating that pinned vortices do not contribute to magnetic noise in the GHz frequency range.
@@
$T_{2}$ noise spectroscopy of slow vortex dynamics — At low temperatures, one enters deep into the superconducting phase and the vortex motion freezes out. Nevertheless, with $H>H_{c1}$, BSCCO can exhibit complex vortex dynamics, such as thermally assisted flux jumps between different pinning centers, or de-pinning of vortices to exit the superconducting sample, potentially leading to a revival of magnetic noise. However, our data for $\Gamma_{1}(T)$ in the low temperature regime exhibits no additional features at finite $H$. One possibility is that, unlike the metal-superconducting transition, fluctuations arising from vortex dynamics at low temperatures occur at a much lower frequency range, and therefore eludes detection by $T_{1}$ spectroscopy that is only sensitive to the noise spectrum in the GHz range.
@@
To this end, we perform NV decoherence ($T_{2}$) noise spectroscopy to capture the magnetic noise spectrum in the MHz frequency range machado2023quantum ; xue2024signatures ; ziffer2024quantum ; li2024critical . Specifically, after preparing the NV center into a quantum superposition state, $(|m_{s}=0\rangle+|m_{s}=-1\rangle)/\sqrt{2}$, we apply an advanced dynamical decoupling sequence (XY-8, see Supplementary Information) to measure the decoherence rate induced by BSCCO, $\Gamma_{2}(T)$, as a function of both temperature and bias external field $H$ [Fig. 4(b)]. Remarkably, $\Gamma_{2}(T)$ exhibits a broad noise spectrum with several prominent features. (i) The onset temperature for $\Gamma_{2}(T)$ progressively decreases with increasing field $H$, and does not coincide with the peak of $\Gamma_{1}$ near criticality. (ii) For a fixed $H$, $\Gamma_{2}(T)$ reaches its maximum value at a much lower temperature $T^{*}$ than $\Gamma_{1}(T)$; however, $H$ is larger than the independently measured $H_{c1}$ of the sample at $T^{*}$ (see Methods), indicating that vortices are still present in the sample. (iii) The maximum value of $\Gamma_{2}(T)$ is nearly independent of the external field $H$.
@@
To understand this behavior, we begin by noting that the decoherence time $T_{2}$ on and off BSCCO is nearly identical in both the metallic phase and near criticality [inset of Fig. 4(b)]. This observation indicates that $T_{2}$ noise spectroscopy is less sensitive than $T_{1}$ in detecting both Johnson-Nyquist noise and critical current fluctuations. This difference in sensitivity can be attributed to the intrinsic timescale gap between $T_{1}$ and $T_{2}$. Specifically, the NV $T_{1}$ is around three orders of magnitude larger than $T_{2}$, offering a much longer time window to accumulate the noise signals (see Methods). Consequently, a significantly stronger noise amplitude from BSCCO is required to induce a measurable $T_{2}$ in NV sensors.
@@
Furthermore, $T_{2}$ measured across a line-cut intersecting a pinned vortex show no position dependence, indicating that frozen vortices also do not contribute to magnetic noise in the MHz frequency range. Combining these observations, we deduce that $T_{2}$ is most sensitive to vortex motion within the vortex solid phase MFisher89 ; fisherfisherhuse . The substantial noise magnitude suggests that it likely originates from thermal fluctuations, which can cause positional jumps of fluxes or flux bundles between different pinning centers, resulting in sudden large changes in the local magnetic field. These dynamics are relatively slow, extending only up to the MHz range. Thus, they are not reflected in $\Gamma_{1}$, but can be directly captured by $\Gamma_{2}$ [Fig. 4(c)]. Only at very low temperatures, approximately corresponding to the Meissner phase, does such flux motion begin to freeze out, leading to a decrease in $\Gamma_{2}(T)$ below $40$ K.
@@
Summary and Outlook — For the first time, we demonstrated the use of NV centers as nano-scale quantum sensors to probe three distinct aspects of low-energy dynamics in high-$T_{c}$ superconductors: gapless quasiparticle excitations at low temperatures, critical superconducting fluctuations near $T_{c}$, and vortex motion in an intermediate temperature range in presence of an external magnetic field. The nanoscale proximity and high sensitivity of our NV sensors, combined with their operational frequency range spanning three decades (from MHz to GHz) and sub-micron spatial resolution, enable the measurement of local dynamics beyond the capabilities of conventional techniques [Fig. 1(a)].
@@
Looking forward, our work opens the door to several intriguing directions. First, while our current studies involve NV center ensembles with a spread of sensor-sample distances, an immediate future direction is to precisely adjust the distance $d$ between the quantum sensor and the target material. This can be achieved by using NVs on a scanning tip maletinsky2012robust ; degen2008scanning ; pelliccione2016scanned , or by creating spin sensors in a given layer of Van der Waals materials such as hexagonal boron nitride gottscholl2020initialization ; gong2023coherent ; stern2022room ; gong2023isotope . By tuning the distance $d$, we can adjust the momentum filter function to enable controlled probing over a variety of length scales (from nanometers to micrometers)  machado2023quantum , which is crucial for imaging inhomogeneous dynamics in superconductors bhattacharyya2023imaging ; Dailledouze2025 .
@@
Moreover, our measurement of the critical slowing down time-scale $\tau_{0}$ places strong constraints on theoretical models of the transition and provides new opportunities to study the mechanisms for order-parameter relaxation near criticality. In fact, a reasonable agreement with the weak coupling result $\tau_{0}=\tau_{\rm GL}$ on the metallic side larkin2005theory , but a stronger deviation on the superconducting side schmid1966time , indicates the necessity for further microscopic investigations of pair dynamics in unconventional superconductors.
@@
Finally, our technique is not limited to BSCCO, and can be readily extended to a variety of superconducting materials. For instance, certain superconductors such as hydrides pickard2020superconducting ; drozdov2015conventional and nickelates sun2023signatures ; wen2024probing undergo superconducting transitions under high pressure. Thus, moving beyond temperature-driven transitions, the integration of NV centers into high-pressure diamond anvil cells offers a pathway to study how pressure influences critical fluctuations and vortex dynamics hsieh2019imaging ; lesik2019magnetic ; yip2019measuring ; bhattacharyya2023imaging ; Dailledouze2025 .
@@
Acknowledgements: We gratefully acknowledge discussions with Pavel Dolgirev, Chris Laumann, Peng Cai, Panyu Hou, Srinivas Mandyam, and Zhipan Wang. This work is supported by NSF NRT LinQ 2152221, NSF ExpandQISE 2328837, and the Center for Quantum Leaps at Washington University. O. K. D. and F. M. acknowledge support from the NSF through a grant for ITAMP at Harvard University. The work at Brookhaven National Laboratory is supported by the US Department of Energy, office of Basic Energy Sciences, contract No. DOE-sc0012704. E. A. H. and X. D. acknowledge support by the Gordon and Betty Moore Foundation, grant DOI 10.37807/gbmf11560. S. R., Q. X., and Z. R. acknowledge support under National Science Foundation (NSF) Division of Materials Research Award DMR-2236528.

### v2

@@ Abstract
To quantitatively analyze the observed magnetic noise, we extract the BSCCO-induced decay rate, $\Gamma_{1}(T)$, by subtracting the intrinsic relaxation rate — measured from NV centers located far from the BSCCO sample — from the total relaxation rate of NV centers situated beneath the BSCCO [Fig. 3(d)]. At high temperatures, $T>T_{c}$, Johnson-Nyquist noise in the metallic phase determines the strength of the magnetic noise [43, 1, 35]. At very low temperatures, $T\ll T_{c}$, the magnetic noise is significantly suppressed. Crucially, the form of this suppression offers insights into the nature of the superconducting pairing symmetry. While in conventional s-wave superconductors $\Gamma_{1}(T)$ is exponentially suppressed at low temperatures due to a fully gapped Bogoliubov quasiparticle spectrum  [5, 4, 11, 18, 38, 49], the presence of nodal quasiparticles in d-wave layered superconductors implies a distinct power-law suppression of $\Gamma_{1}(T)\sim T^{2}$ [18, 11]. Indeed, this expectation is consistent with our observations [Fig. 3(d), inset], corroborating the nodal d-wave nature of superconductivity in BSCCO [37].
@@
The most striking feature of Fig. 3(d) is the sharp and symmetric peak in $\Gamma_{1}(T)$ near the critical temperatures $T_{c}$. This peak arises from strong fluctuations in the superconducting order parameter, resulting in enhanced current fluctuations that act as source of magnetic noise at the metal-superconductor phase transition. Crucially, this enhancement is observed across the entire BSCCO sample [Fig. 3(c)]. Interestingly, near $T_{c}$, $\Gamma_{1}(T)$ begins to show marked deviations from mean-field expectations for d-wave superconductors. In particular, mean-field theory predicts a divergence of $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$ (up to logarithmic corrections) [18]. By contrast, the experimentally measured $\Gamma_{1}(T)$ at criticality is distinctly more singular, with $x\approx 1$ [Fig. 3(e)]. This deviation arises because BCS theory neglects both amplitude and phase fluctuations of the superconducting order parameter $\psi$, which play a significant role near the critical point [21, 46, 53]. These effects are particularly important for thin-film superconductors where fluctuations are expected to be stronger [70]. Consequently, a quantitative characterization of the divergent time-scale associated with dissipative dynamics of the superconducting order parameter near $T_{c}$ remains an important open question, which we turn to next.
@@
To properly account for low-energy critical fluctuations, we consider the following phenomenological Langevin dynamics of the superconducting order parameter $\psi(\mathbf{r},t)$:
@@
where $F[\psi]$ is the Ginzburg-Landau free energy functional [44, 14, 66, 70], $\eta$ is local white noise that captures the effect of coarse-graining, and $\gamma$ sets the rate of relaxation towards equilibrium [33]. Crucially, by setting $r(T)\propto T-T_{c}$ near the critical point, we can analyze fluctuations of the order parameter around the minima of free energy associated with both phases: $\langle\psi\rangle=0$ for $T>T_{c}$ (metallic) and $\langle\psi\rangle=\psi_{0}$ for $T<T_{c}$ (superconducting). These order parameter fluctuations manifest themselves as current fluctuations in the BSCCO sample which induce magnetic noise at the NV’s frequency $D_{\mathrm{gs}}$, enhancing its relaxation rate $\Gamma_{1}(T)$. This low-frequency fluctuation enhancement is directly connected to the phenomena of critical slowing down near the phase transition; the noise spectral density at low frequencies scales with the scattering timescale which diverges as one approaches the critical point. For weak-coupling superconductors, this divergent timescale is quantified by $\tau_{\rm GL}=8\hbar/(k_{B}|T-T_{c}|)$ near $T_{c}$ [46, 65].
We find that our theoretical model (see Methods) can faithfully reproduce the divergence of $\Gamma_{1}(T)$ on both sides of the transition [Fig. 3(e)]. At the same time, our approach enables us to quantitatively extract the value of the scattering timescale $\tau^{\text{fit}}$ by fitting our model to the experimental data on both the metallic [M] and superconducting [SC] sides of the transition. Remarkably, the extracted value for the scattering timescale, given by $\tau^{\rm fit}_{\rm M[SC]}/\tau_{\rm GL}\approx 1.2[0.2]$, are comparable to the analytical weak-coupling results, $\tau^{\rm wc}_{\rm M[SC]}/\tau_{\rm GL}=1[0.5]$ [46, 65, 39]. Understanding the broad agreement between the weak-coupling prediction and the expected strong-coupling nature of BSCCO superconductivity remains an open question [47].
The agreement between our theoretical model and the experimental data suggests that the critical behavior is properly captured by our Langevin dynamics approach. Indeed, the resulting value of the associated critical exponents, $\nu=1/2$ and $z=2$ [33], directly reproduce the aforementioned divergence of the scattering timescale $\tau$ via the scaling behavior of the correlation length $\xi$: $\tau\propto\xi^{z}\propto|T-T_{c}|^{-z\nu}$.
@@
Figure 4: In-field criticality and vortex dynamics. (a) BSCCO-induced relaxation rate, $\Gamma_{1}(T)$ (normalized), as a function of temperature at $H=0,40,200$ G. Inset: the measured peak magnitude of $\Gamma_{1}(T)$ as a function of applied field, which is consistent with a linear dependence on the field strength $H$ (dashed line). (b) BSCCO-induced decoherence rate, $\Gamma_{2}(T)$, as a function of temperature at $H=55,120,200$ G. Inset: The measured $T^{\mathrm{XY8}}_{2}$ decoherence timescale as a function of temperature for NVs under and away from BSCCO at $H=200$ G. (c) Schematic of the noise spectrum arising from different dynamical processes in BSCCO in the context of our two probing techniques, NV relaxation ($T_{1}$) and decoherence ($T_{2}$) noise spectroscopy. While less sensitive than $T_{1}$ spectroscopy, decoherence spectroscopy is able to access fluctuations at much lower frequencies and observe vortex solid dynamics. (d) Left: a trapped vortex imaged in BSCCO under $H=40$ G at 50 K. Right: the measured relaxation time $T_{1}$ and decoherence time $T^{\mathrm{XY8}}_{2}$ of NV centers at different spatial positions across the vortex.
@@
In-field criticality and vortex dynamics — Next we turn on the applied external field and explore the in-field critical fluctuations at $H=40~{}$G and $H=200~{}$G. Intriguingly, the measured $\Gamma_{1}(T)$ differs from the zero-field case in two key ways [Fig. 4(a)]: (i) The peak position of the noise is slightly shifted towards lower temperatures. (ii) The peak becomes broad and highly asymmetric; while the noise signal remains sharp when approaching criticality from the metallic side, it increases much more gradually and smoothly from the superconducting side. The former is expected, as a perpendicular field hinders phase coherence and lowers the critical temperature $T_{c}$. The latter can be attributed to two different factors. On the one hand, an external magnetic field broadens the temperature range where critical fluctuations are important as described by two-dimensional XY critical dynamics [7, 29, 13]. On the other hand, below $T_{c}$, an intermediate vortex liquid phase is expected to form where the local motion of the superconducting vortices (which trap magnetic flux quanta) can generate an additional fluctuating magnetic field at the NV [9, 23, 58].
@@
Further evidence for the latter can be gleaned from the scaling of the peak magnitude of $\Gamma_{1}$ with the magnetic field $H$ [inset, Fig. 4(a)], where the density of vortices $n_{v}$ increases with increasing field strength. Within a simple model of diffusive vortex motion (see Methods), the local magnetic noise, measured via $\Gamma_{1}(H)$, scales as $D_{v}^{-1}$ where $D_{v}\sim v^{2}\tau_{v}$ is the vortex diffusion constant. The vortex scattering time $\tau_{v}$ is expected to be inversely proportional to both the vortex density $n_{v}$, and the mean-square speed $\langle v^{2}\rangle\sim T$ is primarily determined by the temperature of the sample. Consequently, the measured magnetic noise scales as $\Gamma_{1}(H)\propto n_{v}\propto H$, in agreement with our observations [inset, Fig. 4(a)]. As the temperature further decreases, the vortices either freeze into a lattice or become pinned by defects, resulting in a decrease of $\Gamma_{1}(T)$. This conclusion is consistent with our measurement of noise over a line-cut on a large pinned vortex at low temperatures shown in Fig. 4(d) — $T_{1}$ does not change appreciably across the vortex, indicating that pinned vortices do not exhibit dynamics in the GHz frequency range.
@@
$T_{2}$ noise spectroscopy of slow vortex dynamics — At sufficiently low temperatures, the system is deep in the superconducting phase and the vortex liquid is expected to transition into a solid state. Nevertheless, with $H>H_{c1}$, BSCCO can still exhibit complex vortex dynamics, such as thermally assisted flux jumps between different pinning centers, or the de-pinning of vortices to exit the superconducting sample, leading to fluctuating magnetic signals from different physical processes. However, the measured $\Gamma_{1}(T)$ at low temperature regime does not display any features, suggesting that these processes operate at even lower frequency, and therefore elude detection by $T_{1}$ spectroscopy sensitive only to GHz-range fluctuations [Fig. 4(c)].
@@
To investigate these dynamics, we perform NV decoherence ($T_{2}$) noise spectroscopy to capture the magnetic noise spectrum in the MHz frequency range [52, 73, 76, 50]. Specifically, after preparing the NV center into a quantum superposition state, $(|m_{s}=0\rangle+|m_{s}=-1\rangle)/\sqrt{2}$, we apply a dynamical decoupling sequence (XY-8, see Supplementary Information) to measure the decoherence rate induced by BSCCO, $\Gamma_{2}(T)$, as a function of both temperature and applied field $H$ [Fig. 4(b)]. Remarkably, $\Gamma_{2}(T)$ exhibits a broad noise spectrum with several prominent features. (i) With increasing field $H$, there is a reduction of the onset temperature for $\Gamma_{2}$-type noise and this onset occurs for temperatures below the peak position of $\Gamma_{1}(T)$. (ii) For different $H$, $\Gamma_{2}(T)$ reaches its maximum value at a much lower temperature $T^{*}$ than $\Gamma_{1}(T)$; however, the applied field $H$ remains larger than the independently measured $H_{c1}$ of the sample at $T^{*}$ (see Methods), indicating that vortices are still present in the sample. (iii) The maximum value of $\Gamma_{2}(T)$ is insensitive to the strength of the external field $H$.
@@
To understand these observations, we must rule some of the previously studied mechanisms. In particular, we begin by noting that the decoherence time $T_{2}$ on and off BSCCO is nearly identical in both the metallic phase and near criticality [inset of Fig. 4(b)]. This observation indicates that $T_{2}$ noise spectroscopy is less sensitive than $T_{1}$ in detecting both Johnson-Nyquist noise and critical current fluctuations. This difference in sensitivity can be attributed to the separation between the intrinsic depolarization and dephasing timescales of the NV. More specifically, the NV’s $T_{1}$ is around three orders of magnitude larger than its $T_{2}$; a significantly stronger noise amplitude is required to induce a measurable change of NV $T_{2}$ (Methods).
@@
At the same time, $T_{2}$ measured across a line-cut intersecting a strongly pinned vortex shows no position dependence, indicating that static vortices also do not contribute to magnetic noise in the MHz frequency range. This suggests that $T_{2}$ is most sensitive to vortex motion within the vortex solid phase [24, 23]. More specifically, the noise magnitude implies that it likely originates from thermal fluctuations, which can cause positional jumps of fluxes or flux bundles between different weak pinning centers, resulting in sudden large changes in the local magnetic field. These dynamics are relatively slow, extending only up to the MHz range. Thus, they are not reflected in $\Gamma_{1}$, but can be directly captured by $\Gamma_{2}$ [Fig. 4(c)]. Only at very low temperatures, approximately corresponding to the Meissner phase, does such flux motion begin to freeze out, leading to a decrease in $\Gamma_{2}(T)$ below $40$ K.
@@
Outlook — For the first time, we demonstrate the use of NV centers as nano-scale quantum sensors of superconducting dynamics. The nano-scale proximity and high sensitivity of our NV sensors, combined with their operational frequency range spanning three decades (from MHz to GHz) and sub-micron spatial resolution, enable the measurement of a wide swath of dynamical phenomena beyond the capabilities of conventional techniques [Fig. 1(a)].
@@
Looking forward, our work opens the door to several intriguing directions. First, while our current studies involve ensemble NV centers with a spread of sensor-sample distances, an immediate future direction is to precisely adjust the distance $d$ between the quantum sensor and the target material. This can be achieved by using single NV center on a scanning tip [54, 17, 61], or by creating spin sensors in a given layer of Van der Waals materials such as hexagonal boron nitride [28, 27, 68, 26]. By tuning the distance $d$, we can adjust the momentum filter function to enable controlled probing over a variety of length scales (from nanometers to micrometers)  [52], which is crucial for imaging inhomogeneities in superconductors [8, 15].
@@
Moreover, our measurement of the critical slowing down time-scale $\tau$ places strong constraints on theoretical models of the transition and provides new opportunities to study the mechanisms for order-parameter relaxation near criticality. In fact, a reasonable agreement with the weak coupling result $\tau=\tau_{\rm GL}$ on the metallic side [46], but a stronger deviation on the superconducting side [65], indicates the necessity for further microscopic investigations of critical dynamics in unconventional superconductors.
@@
Finally, our technique is not limited to BSCCO, and can be readily extended to a variety of superconducting materials. For instance, certain superconductors such as hydrides [62, 19] and nickelates [69, 72] undergo superconducting transitions under high pressure. Thus, moving beyond temperature-driven transitions, the integration of NV centers into high-pressure diamond anvil cells offers a pathway to study how pressure influences critical fluctuations and vortex dynamics [35, 48, 74, 8, 15].
@@
Acknowledgements: We gratefully acknowledge discussions with Pavel Dolgirev, Chris Laumann, Paul Ching-Wu Chu, Liangzi Deng, Andrew Mounce, Ania Jayich, Peng Cai, Panyu Hou, Srinivas Mandyam, Zhipan Wang. This work is supported by NSF NRT LinQ 2152221, NSF ExpandQISE 2328837, and the Center for Quantum Leaps at Washington University. O. K. D. and F. M. acknowledge support from the NSF through a grant for ITAMP at Harvard University. The work at Brookhaven National Laboratory is supported by the US Department of Energy, office of Basic Energy Sciences, contract No. DOE-sc0012704. E. A. H. and X. D. acknowledge support by the Gordon and Betty Moore Foundation, grant DOI 10.37807/gbmf11560. S. R., Q. X., and Z. R. acknowledge support under National Science Foundation (NSF) Division of Materials Research Award DMR-2236528.